# OriStitch: A Machine Embroidery Workflow to Turn Existing Fabrics into Self-Folding 3D Textiles

### v1

@@ 4.2.3. Dissolving support threads
As illustrated in Figure 3b the heat actuated thread is held in place with a water soluble bobbin thread4. During embroidery, something needs to hold the actuated thread in place, which is where this specific bobbin thread comes in. As the figure shows, the bobbin thread is very low in tension to not pull the shrinking thread through the fabric, (as opposed to the other grey bobbin threads for lock and channel stitches which are pulled in place). When exposed to water, these bobbin threads dissolve allowing the heat actuated thread to shrink unconstrained.
@@ 4.3. Heat-actuating the hinges into their final shape
To actuate the pattern, we propose dipping the sheet in boiling water, if need by followed by post-processing with a heat gun as shown in Figure 9. The combination of hot water and heat gut treatment enables a relatively fast and reliable actuation with minimal manual labour. The hot water pre-creases all hinges to ±70% of their folded state in a matter of seconds. Models consisting of a small number of triangles completely fold in boiling water because the lever each fold pulls is shorter. The focused distribution of heat with the heat gun makes that each hinge closes one at a time instead of all at once (in water). The time needed for post-processing actuation mainly depends on the number of wide hinges (type 3 and type 4). For example, the hemisphere (55 hinges with 13 wide hinges) in figure 17 was finished in 2 minutes while around 6 minutes were used to fully closed the hinges of the vase cover (144 hinges with 49 wide hinges).
We attempted to use a convection oven as well, which did actuate all hinges in parallel, however simultaneous closing caused hinges to pull on one another at the same time, and the high temperature made the actuated threads fragile, by actuating sequentially, we circumvent that problem albeit a bigger manual effort. Some hinges require more tension than others because of their width and neighboring faces, therefore the manual process allows actuating each hinge until it fully closes. The shrinking conditions depend on various factors including the temperature, the amount of air flow, the distance between the hot air and the threads, and the duration of treatment. Therefore, instead of proving a specific temperature, we recommend using a simple gauge shown in Figure 10 to calibrate the heating conditions for the used heating mechanism and materials.
@@ 4.4. Computational tools to support the OriStitch workflow
Finally, to give users a chance to benefit from our hinge design and produce interesting 3D structures, we provide computational tools to (1) convert a target 3D mesh into machine-instructions and (2) support multi-hooping, a process that allows fabricating larger sheets on small embroidery machines.
@@ 4.4.1. Converting 3D meshes to machine instructions
To invoke OriStitch, users load a target shape as triangle mesh (in the commonly used OBJ format) in the automatic (headless) conversion tool. This tool converts a given mesh into a 2D configuration of faces, built upon the Origamizer (Tachi, 2010) algorithm, which achieves a similar objective in the context of paper folding. The algorithm transforms a disk-equivalent mesh into 2D by calculating each face’s rigid transformation and introducing connecting creases. These creases tuck away excess material either at the edges or vertices. Origamizer obtains such mapping by minimizing an energy function that penalizes self-intersecting faces and invalid creases.
To make this work for embroidery structures of OriStitch, we extend the Origamizer algorithm. Origamizer only avoids self-intersections via a soft constraint penalty term, leading to failures (Figure 11). In paper folding, the excess material at an intersection tucks away just fine, but it leaves insufficient space to insert our embroidered hinges, OriStitch therefore aims to merge intersecting faces. These intersections fall into three categories based on the amount of overlap between the faces; (a) if the edges fully overlap, OriStitch simply merges them into a larger face without inserting a hinge, (b) if the angle between the intersecting edges is small (¡10∘), OriStitch merges the faces while cutting away the non-overlapping geometry, and (c) if the angle exceeds that threshold, OriStitch inserts a hinge between the two faces and separates them. Another fabrication-specific deviation of the Origamizer algorithm is that OriStitch removes material between the faces and hinges, this relaxes optimization constraints to reduce the chance of intersections.
@@
4.4.2. multi-hooping support
Embroidery machines are limited in the size of structures they can fabricate by the size of the hoop that holds the material in place. Multi-hooping is a commonly used technique to extend this range where users fabricate one patch, then release and affix the hoop on an adjacent path of material to extend the fabrication area. To this end we built a custom post-processing step integrated in Inkscape next to common embroidery software Ink/Stitch666https://www.inkstitch.org to feed into pipelines users are familiar with.
@@
This tool strategically subdivides the stitch pattern based on the machine-specific hoop size as shown in Figure 12. Unlike typical multi-hooping techniques, our technique takes into account the specific sequence of fabrication steps for our hinges to work, and the division within hinges, Figure 12b shows how the pink (hoop 2) face is laid out first, and then in green (hoop 3) the active thread is looped in and the lock and channel stitches are added. When users export their fabrication files, the resulting stitch pattern is subdivided into separate hoops to fabricate.
@@ 5. Evaluation
To evaluate our contributions we conducted two technical evaluations. First, we evaluated the success rate of our shape conversion tool, and secondly we evaluate the our fabrication workflow and hinge behavior on a range of different materials.
5.1. OriStitch automatically converted 23/28 mesh models to stitch patterns
To evaluate the mesh conversion, we used OriStitch to replicate mesh models found in related research papers. We measured the success rate, the types of hinges used, and the number of merged triangles. We defined success as meshes which produced a full stitch pattern and had no intersections between faces/hinges.
@@ 5.1.1. Models
To create a fair benchmark for our technique, we replicate models found in four related papers which also develop 3D structures from flat sheets (Guseinov et al., 2017; Narumi et al., 2023; Panetta et al., 2021; Jourdan et al., 2023). We furthermore add six that would not be supported by those techniques as they produce 2.5D textures on materials or construct a more advanced topology from separate sheets stitched together.
@@ 5.1.2. Pre-processing
As mentioned before, OriStitch has a minimum resolution of each face edge. To accommodate for this, we re-meshed each model to regular triangles, for very small models we scaled the model up to overcome these limitations. We furthermore verify that vanilla Origamizer did indeed succeed for all models, so all findings are related to the OriStitch extension.
@@ 5.1.3. Results
Figure 14 shows the models and resulting stitch patterns generated by OriStitch. 23/28 of the models will correctly fabricate. For the remaining 5/28 models there were intersections which OriStitch could not resolve.
Table 1 furthermore shows the variety of types of hinges used by OriStitch to make these models fabricate (type 1,2,3,4) and the minimal edge length of the faces. The models which required scaling during pre-processing all have in common that the minimum edge length matches the resolution of OriStitch.
| model | faces | inters. | hinge types |
| Grid | 443 | 0 | 3,107,459,77 |
| Heart | 590 | 1 | 1,33,769,53 |
| Teapot | 4455 | 3 | 3,77,1819,1720 |
| Letter | 1880 | 5 | 2,25,1253,1481 |
| Torus | 803; 2 cuts | 1 | 12,199,877,62 |
| Tessellation | 24 | 0 | 0,14,6,8 |
| Hat | 479 | 3 | 1,136,255,300 |
| Car | 941 | 4 | 8,261,676,432 |
| Half Sphere | 100 | 0 | 0,73,44,25 |
| Hyperboloid | 82 | 4 | 4,21,32,48 |
| Turtle | 312 | 5 | 9,55,212,152 |
| Bump Cap | 340 | 0 | 1,135,225,132 |
| Miura Fold | 72 | 1 | 5,35,6,47 |
| Waterbomb | 60 | 6 | 6,15,38,11 |
| Bauhaus paraboloid | 140; 1 cut | 3 | 7,62,52,58 |
| Yoshimura Pattern | 28 | 0 | 0,3,47,23 |
| Mountain | 112 | 0 | 0,23,96,36 |
| Saddle | 98 | 2 | 0,15,71,47 |
| Annulus | 2542; 1 cut | 0 | 0,121,2134,1463 |
| Two Bumps | 1435 | 1 | 5,111,1216,768 |
| Cashew | 658 | 0 | 5,172,455,308 |
| Lilium | 185 | 1 | 3,40,123,91 |
| Cast | 250;1 cut | 0 | 0,46,172,134 |
@@ 5.1.4. Discussion
In total, OriStitch successfully converted 23/28 meshes into fabricatable stitch patterns.
As shown in Figure 13 we encountered all forms of merging in our models. Especially the green models with relatively small features and sharp angles exhibited a frequent need for merging of face. OriStitch does warn users of the number of merged faces as depending on the material used, it may impact the actuation process. We also found that in cases where the mean curvature is very low (local maximum or minimum) but the Gaussian curvature is not (like the pink saddle shape or the orange hat), which implies this shape is locally not developable, Origamizer places faces very far away from one another depending on the steepness of the curvature. This results in a cycle of type 4 hinges in OriStitch. Technically this can be fabricated, but because there is a lot of crimping that has to happen in a small area, it may produce an excess of tucked away material and thus a reduction of hinge tightness.
Closer inspection into the five failing models shows that in these cases there were several intersections in adjacent faces, this is caused by heavy fluctuation of curvature, Figure 15 shows these models and highlights the vertex of maximum curvature in each. OriStitch attempts to accommodate by merging intersecting faces, but in doing so changes the hinges to neighboring faces resulting in new intersections. They contain regions that are locally hard to develop, thus producing faces that are spaced out too far (while having the merging issues indicated before). To avoid this in the future, we plan to explore the use of Gauss thinning (Binninger et al., 2021), resulting in better local developability of our meshes. For the time being, OriStitch detects these cases and throws an error message for users before they aim to fabricate resulting stitch patterns.
@@
One of the core advantages of embroidery is that the additive textile manufacturing method allows it to work on a range of off-the-shelf textiles. We evaluated whether and how OriStitch’ hinges are influenced by material variations. We applied a simple stitch pattern consisting to each sample.
@@ 5.2.1. Stitch pattern
We used a basic stitch pattern as shown in Figure 16 consisting of four fully connected faces (one being at the minimum resolution) and with three different hinge types to validate the full functionality of OriStitch hinges across materials. We also validate our hinges designs with a larger structure, here a hemisphere, consisting of the combination of several hinges.
@@ 5.2.2. Materials
We selected the following materials: neoprene777https://burlapfabric.com/index.php?main_page=product_info&products_id=2933, non-woven (leather)888https://www.amazon.com/dp/B08CF43C7X, felt999https://www.amazon.com/dp/B01GCLS32M , a composite material (natural cork with Polyurethane Mesh on the back)101010https://www.amazon.com/dp/B0B4FNJPDW, and even-woven fabric (Aida 18count)111111https://www.michaels.com/product/loops-threads-aida-cloth-cross-stitch-fabric-295-x-36-18-count-10357725.The thickness of the textile ranges from 0.3mm to 2mm. Table 2 shows an overview of the different samples.
@@ 5.2.4. Discussion
OriStitch hinges worked reliably across material samples. For thicker materials, we do recommend using a thicker heat-actuated thread as the tension in the thread appeared to be the limiting factor for the thicker, less flexible materials. We furthermore saw that the material sample with a low density of fibers poses a risk of failure in the widest hinges. In our evaluation we adjusted by reducing the speed of the machine, but we plan to incorporate a feature in the software to add additional needle points (reducing the bridging of the wide hinge) for more delicate materials.
Finally, when the materials are too soft, they are unable to withstand the tension within faces. Our Neoprene still worked in the end, but some minimal surface stiffness appears to be required for predictable folding.

### v2

@@ 4.2.3. Dissolving support threads
As illustrated in Figure 3b, the heat-actuated thread is held in place with a water-soluble bobbin thread4. During embroidery, something needs to hold the actuated thread in place, which is where this specific bobbin thread comes in. As the figure shows, the bobbin thread is very low in tension so as not to pull the shrinking thread through the fabric (as opposed to the other grey bobbin threads for lock and channel stitches, which are pulled in place). When exposed to water, these bobbin threads dissolve, allowing the heat-actuated thread to shrink unconstrained.
@@ 4.3. Heat-actuating the hinges into their final shape
To actuate the pattern, we propose dipping the sheet in boiling water, if needed, followed by post-processing with a heat gun as shown in Figure 9. The combination of hot water and heat gut treatment enables a relatively fast and reliable actuation with minimal manual labor. The hot water pre-creases all hinges to ±70% of their folded state in a matter of seconds. Models consisting of a small number of triangles completely fold in boiling water because the lever each fold pulls is shorter. The focused distribution of heat with the heat gun ensures that each hinge closes one at a time instead of all at once (in water).
We attempted to use a convection oven as well, which did actuate all hinges in parallel, however simultaneous closing caused hinges to pull on one another at the same time, and the high temperature made the actuated threads fragile, by actuating sequentially, we circumvent that problem albeit a bigger manual effort. Some hinges require more tension than others because of their width and neighboring faces, therefore the manual process allows actuating each hinge until it fully closes. The shrinking conditions depend on various factors, including the temperature, the amount of air flow, the distance between the hot air and the threads, and the duration of treatment. Therefore, instead of proving a specific temperature, we recommend using a simple gauge shown in Figure 10 to calibrate the heating conditions for the used heating mechanism and materials.
@@ 4.4. Computational tools to support the OriStitch workflow
Finally, to give users a chance to benefit from the Oristitch hinge design to produce interesting 3D structures, we provide computational tools to (1) convert a target 3D mesh into machine instructions and (2) support multi-hooping, a process that allows fabricating larger sheets on small embroidery machines.
@@ 4.4.1. Converting 3D meshes to machine instructions
Oristitch’s interface is a simple GUI based on Tachi’s Origamizer (Tachi, 2010). This tool converts a given mesh into a 2D configuration of faces, built upon the Origamizer (Tachi, 2010) algorithm, which achieves a similar objective in the context of paper folding. The algorithm transforms a disk-equivalent mesh into 2D by calculating each face’s rigid transformation and introducing connecting creases. These creases tuck away excess material either at the edges or vertices. The Oristitch interface (shown in Figure 11) allows users to tune the resulting crease patterns interactively, includes machine-specific constraints, subdivides the pattern for multi-hooping, and exports fabrication-ready machine instructions for laser cutters and embroidery machines.
Oristitch is an extension of the two-phase mapping algorithm for origamizing polyhedral surfaces, as introduced in Origamizer (Tachi, 2010). We preserve the original framework’s core strength in ensuring geometric foldability and focus our contribution on the Translational Mapping phase, where the final positions of the unfolded polygons are determined.
The original algorithm in this phase uses an iterative optimization process to find a valid, non-overlapping layout. Its primary objective is to minimize a penalty energy, , derived from a set of geometric and topological constraints. We augment this process by introducing a new, tunable penalty term that allows a designer to enforce a user-defined minimum separation distance, a critical constraint for physical fabrication as machine and material properties enforce their own minimum separation for successful fabrication, and users may have aesthetic preference to further increase the distance at the trade-off of a larger sheet to fabricate. Oristitch’ user interface shows users the size, and number of each hinge type, which updates interactively as the user manipulates the stitch pattern, as shown in Figure 11.
Oristitch accomplishes this by modifying the objective function that the optimizer minimizes. The total penalty energy, , is now a weighted sum of the original geometric penalty and our new distance penalty:
Here, represents the original penalties from Tachi’s method. The new term, , enforces our minimum distance rule, , and is a user-defined weight to control its influence. The penalty is formulated as a one-sided quadratic function; it is zero if the separation between two parts is greater than or equal to and grows quadratically as the separation distance drops below this threshold. This creates a strong, stable repulsive force that enforces the spacing constraint.
Crucially, this penalty is applied to the algorithm’s internal proxy distance metric. Oristitch leverages this existing proxy because, while not strictly identical to the Euclidean distance, it is a smooth and computationally stable function of the layout parameters. The gradient of this proxy metric is continuous, which is essential for the stability of the numerical solver. Using this well-behaved metric ensures that the new penalty integrates seamlessly into the original gradient-based optimization framework, preserving the robustness and convergence properties of the foundational algorithm, while respecting fabrication constraints and user preferences.
@@
4.4.2. Multi-Hooping
A single embroidery hoop may not suffice to contain the entire unfolded layout. While large industrial machines should be able to cover most basic stitch pattens, there are limitations for smaller (household) machines, or for very large patterns. A common strategy in embroidery to address this is called ”multi-hooping”. Operators effectively hoop a large sheet of fabric into a smaller hoop, they stitch the first pattern, then ”rehoop” by placing the hoop on a new part of the sheet, typically slightly overlapping with the previously embroidered pattern. The graph nature of Oristitch folding patterns (each face connects to the others by the active thread) makes it non-trivial to naively divide the total design into individual hoops.
@@ 4.4.2. Multi-Hooping
Oristitch therefore strategically subdivides the stitch pattern based on the machine-specific hoop size as shown in Figure 12. Unlike typical multi-hooping techniques, our technique takes into account the specific sequence of fabrication steps for our hinges to work, and the division within hinges, Figure 12b shows how the pink (hoop 2) face is laid out first, and then in green (hoop 3) the active thread is looped in and the lock and channel stitches are added. When users export their fabrication files, the resulting stitch pattern is subdivided into separate hoops to fabricate.
To subdivide the stitch pattern, Oristitch computes the minimal number of vertical or horizontal hoop placements required to fully cover all hinges and their corresponding faces (we call ”hinge-pairs”).
Boundary Extraction.
Each valid gluing edge in the polyhedral model defines a hinge-pair. Instead of computing the precise shape boundary, we use the axis-aligned bounding box (AABB) of these vertices to represent its extent. That is, for each unit we store its , , , . This representation will not lose precision because the we assume axis-aligned hoops (slightly more compact hooping may be achieved by allowing for rotation, but as we will see later, comes at a significant computational cost).
Problem Formulation.
Each hinge-pair is bounded by an axis-aligned rectangular bounding box , and the set of all such boxes is denoted by . The objective is to cover all rectangles in using the minimum number of embroidery hoops, each modeled as a rectangle of fixed width and height (defined by users depending on their available hoop size).
This is a set covering problem, known to be NP-hard. Therefore, we propose a practical greedy heuristic that efficiently places hoops to maximize coverage. Because of the fabrication sequence constraints, each face has to be fully contained in a hoop; taking the AABB of hinge-pairs will safeguard that property. The cyclical nature of the stitch pattern (all face edges except for the boundary are connected to other edges) ensures that there is always required overlap between neighboring hoops to fabricate the stitches shared among hinges.
Greedy Multi-hooping algorithm.
The algorithm proceeds iteratively as follows:
Anchor Candidate Extraction. Compute the union of all bounding boxes in the hinge-pair set , where each unit is represented by an axis-aligned rectangle . To identify a discrete set of candidate anchor positions for hoop placement, we collect all distinct - and -coordinates from the bounding box edges:
(14) The Cartesian product defines a discrete set of anchor candidates, where each corresponds to a potential lower-left corner for placing a hoop.
Hoop Simulation. For each , simulate placing a hoop at that anchor with two orientations:
Horizontal: hoop size , bottom-left corner at ;
Vertical: hoop size , bottom-left corner at .
Count how many remaining boxes in are fully contained within each simulated hoop.
Greedy Selection. Select the hoop placement (position and orientation) that covers the largest number of bounding boxes. If no placement covers any remaining box, the algorithm terminates with failure.
Update. Remove all covered boxes from and repeat from step 1 until is empty.
Runtime Analysis.
Let be the number of Hinge-pairs. Each is represented by an axis-aligned bounding box.
In each iteration, we collect at most distinct and values from all bounding boxes, yielding candidate anchor positions. For each position, we evaluate two hoop orientations and check containment against boxes, resulting in work per iteration.
As each iteration removes at least one unit, there are at most iterations. The total worst-case runtime is:
Thus, the algorithm runs in polynomial time.
4.4.3. Export Stitch Pattern
Once the unfolded model is developed and the hoop placement is completed, users export a machine-ready SVG file for fabrication. The generated SVG is structured in a hierarchical grouping format to reflect the embroidery and laser cutting processes.
As shown in Figure 13, at the top level, each group corresponds to either a specific hoop placement or the laser-cut layer. Within each hoop group, subgroups are used to differentiate between various types of stitching patterns, such as mountain folds, valley folds, locking stitches, and others. Furthermore, within each stitching pattern subgroup, the individual paths are sorted using a greedy algorithm. We use a greedy distance-based algorithm that iteratively selects the next nearest stitching path to minimize travel distance and reduce non-stitching jumps during embroidery. This sorting approach helps reduce excessive jumping between stitches and facilitates a more efficient machine operation.
This structured output allows direct integration with embroidery machines and supports both laser cutting and stitching operations. The layering ensures that the fabrication order is preserved and interpretable.
@@ 5. Evaluation
To evaluate our contributions, we conducted two technical evaluations. First, we evaluated the success rate of our shape conversion tool, and secondly, we evaluated our fabrication workflow and hinge behavior on a range of different materials.
5.1. OriStitch automatically converted 26/28 mesh models to stitch patterns
To evaluate the mesh conversion, we used OriStitch to replicate mesh models found in related research papers. We measured the success rate, the types of hinges used, and the size of the overall stitch plan.
@@ 5.1.1. Models
To create a fair benchmark for our technique, we replicate models found in four related papers, which also develop 3D structures from flat sheets (Guseinov et al., 2017; Narumi et al., 2023; Panetta et al., 2021; Jourdan et al., 2023). We furthermore add six that would not be supported by those techniques as they produce 2.5D textures on materials or construct a more advanced topology from separate sheets stitched together.
@@ 5.1.2. Pre-processing
OriStitch has a minimum resolution of each face edge. To accommodate for this, we remeshed each model to regular triangles, for very small models, we scaled the model up to overcome these limitations.
@@ 5.1.3. Results
Figure 14 shows the models and resulting stitch patterns generated by OriStitch. 26/28 of the models will correctly fabricate. The 2 models that did not work failed because the minimum resolution of Oristitch triangles and required re-triangulation at that scale degraded the original models to a degree that they no longer resemble the originals.
Table 1 shows the variety of types of hinges used by OriStitch to make these models fabricate (type 0,1,2,3) and the minimal edge length of the faces. The models which required scaling during pre-processing all have in common that the minimum edge length matches the resolution of OriStitch.
| model | faces | hinge types |
| Grid | 443 | 239,232,126,51 |
| Heart | 590 | 420,236,161,44 |
| Teapot | 2455 | 912,1363,945,400 |
| Letter | 1880 | 766,1304,524,170 |
| Torus | 857; 2 cuts | 475,180,248,351 |
| Tessellation | 24 | 12,14,3,0 |
| Face | 798 | 242,693,208,37 |
| Hat | 479 | 253,388,54,0 |
| Car | 941 | 509,636,221,15 |
| Half Sphere | 100 | 57,66,15,4 |
| Hyperboloid | 82 | 12,81,13,4 |
| Turtle | 312 | 88,172,94,85 |
| Bump Cap | 340 | 169,264,61,0 |
| Mask | 1795 | 303,1170,518,664 |
| Miura Fold | 72 | 27,68,1,0 |
| Waterbomb | 60 | 10,50,6,12 |
| Bauhaus paraboloid | 140; 1 cut | 29,57,23,75 |
| Yoshimura Pattern | 28 | 8,19,1,5 |
| Mountain | 112 | 74,82,0,0 |
| Saddle | 98 | 41,58,32,3 |
| Ronresch | 424 | 212,192,137,65 |
| Annulus | 2542; 1 cut | 557,820,580,1762 |
| Two Bumps | 1435 | 753,688,307,355 |
| Cashew | 658 | 173,288,158,327 |
| Lilium | 185 | 78,152,24,6 |
| Cast | 250;1 cut | 76,147,67,62 |
@@ 5.1.4. Discussion
In total, OriStitch successfully converted 26/28 meshes into fabricatable stitch patterns.
We found that in cases where the mean curvature is very low (local maximum or minimum) but the Gaussian curvature is not (like the pink saddle shape or the orange hat), which implies this shape is locally not developable, Origamizer places faces very far away from one another depending on the steepness of the curvature. This results in a cycle of type 3 hinges in OriStitch. Technically this can be fabricated, but because there is a lot of crimping that has to happen in a small area, it may produce an excess of tucked away material and thus a reduction of hinge tightness.
@@
One of the core advantages of embroidery is that the additive textile manufacturing method allows it to work on a range of off-the-shelf textiles. We evaluated whether and how OriStitch’s hinges are influenced by material variations. We applied a simple stitch pattern to each sample.
@@ 5.2.1. Stitch pattern
We used a basic stitch pattern as shown in Figure 15, consisting of four fully connected faces (one being at the minimum resolution) and with three different hinge types to validate the full functionality of OriStitch hinges across materials. We also validate our hinges designs with a larger structure, here a hemisphere, consisting of the combination of several hinges.
@@ 5.2.2. Materials
We selected the following materials: neoprene666https://burlapfabric.com/index.php?main_page=product_info&products_id=2933, non-woven (leather)777https://www.amazon.com/dp/B08CF43C7X, felt888https://www.amazon.com/dp/B01GCLS32M , a composite material (natural cork with Polyurethane Mesh on the back)999https://www.amazon.com/dp/B0B4FNJPDW, and even-woven fabric (Aida 18count)101010https://www.michaels.com/product/loops-threads-aida-cloth-cross-stitch-fabric-295-x-36-18-count-10357725.The thickness of the textile ranges from 0.3mm to 2mm. Table 2 shows an overview of the different samples.
@@ 5.2.4. Discussion
OriStitch hinges worked reliably across material samples. For thicker materials, we do recommend using a thicker heat-actuated thread as the tension in the thread appeared to be the limiting factor for the thicker, less flexible materials. We furthermore saw that the material sample with a low density of fibers poses a risk of failure in the widest hinges. In our evaluation, we adjusted by reducing the speed of the machine, but we plan to incorporate a feature in the software to add additional needle points (reducing the bridging of the wide hinge) for more delicate materials.
Finally, when the materials are too soft, they are unable to withstand the tension within the faces. Our Neoprene still worked in the end, but some minimal surface stiffness appears to be required for predictable folding.